OpenStudy (anonymous):

If m^6 and m^5= (y^2)/6, then which of the following expresses m in terms of y?

hartnn (hartnn):

what ? both m^6 and m^5 cannot be equal to y^2/6 ...

OpenStudy (anonymous):

If m^6=3/y and m^5= (y^2)/6, then which of the following expresses m in terms of y? Sorry, it's really late right now and I'm tired haha

hartnn (hartnn):

ok, just divide the 2 equations! \(\large \dfrac{m^6}{m^5}=m^{6-5}=m\)

OpenStudy (anonymous):

m is the answer?

hartnn (hartnn):

we just divided left side, still need to divide the right side.. \(\huge \dfrac{\dfrac{3}{y}}{\dfrac{y^2}{6}}=...?\)

hartnn (hartnn):

can u do that division ?

OpenStudy (anonymous):

18/y^3 ?

hartnn (hartnn):

yes, correct :) m = 18/y^3

OpenStudy (anonymous):

Thanks :)

hartnn (hartnn):

welcome ^_^